OpenStudy (anonymous):

in triangle ABC, A=32 degrees, a=8, and b=7. Find C. A. 110.7 degrees B. 20.7 degrees C. 30.4 degrees D. 120.4 degrees

OpenStudy (jack1):

hey @Acenator15 , welcome to OS man, standby will draw a pic: easier then

OpenStudy (jack1):



OpenStudy (jack1):

to solve , either use Sine Rule \[\frac {a}{\sin A} = \frac {b}{\sin B} = \frac {c}{\sin C}\] then triangle rule (sum of all internal angles in a triangle = 180 degrees or use Cos Rule \[a^2 = b^2 + c^2 -(2bc \times \cos A)\]

OpenStudy (jack1):

ur choice...

OpenStudy (jack1):

let me know how u go @Acenator15

OpenStudy (anonymous):

sine

OpenStudy (anonymous):

Are you there

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4fff2465e4b09082c070512f this is the same question may this help you...

OpenStudy (anonymous):

It kind of helped me but they never gave an exact answer so I was stuck

OpenStudy (jack1):

sorry dude, teach a man to fish VS give a man a fish... I'll help u with the solution, not give u the answers

OpenStudy (anonymous):

so what should I do

OpenStudy (jack1):

so check out the picture i drew above... you've got small a , small b and Large A use the sine rule to find Large B (the angle which is opposite side small b) then you'll have angle A and angle B now for every triangle ever, the sum of all the internal angles is 180 degrees so A + B + C = 180 so to find angle C, if you'll know sum of all of the angles = 180 and A = 32 and B = whatever u worked out above so A + B + C = 180 therefore C = 180 - (A + B) : so then you'll know: A, B , C, a, and b ... so use Sine Rule again to find c